OpenStudy (anonymous):

Does this integral converge or diverge, and how? \[\int\limits_{\pi/2}^{\pi}34(cscx)dx\]

OpenStudy (anonymous):

@TuringTest

OpenStudy (anonymous):

it diverges because the graph blows up to infinity if you graph the answer

OpenStudy (anonymous):

That was what I was thinking, until: http://www.wolframalpha.com/input/?i=integral+pi%2F2+to+pi+34cscx @sonofa_nh

OpenStudy (anonymous):

well even that graph...starts at like 48...goes straight..then blows up..so it would diverge

OpenStudy (anonymous):

if it was convergent it would dive down close to zero and stay constant

OpenStudy (anonymous):

i think....it has been a few years since i've taken calc 2....don't really analyze integrals after that class...just do integrals from time to time

OpenStudy (anonymous):

@TuringTest I got stuck when evaluating the ln's

OpenStudy (anonymous):

@dpaInc @Zarkon @ParthKohli @FoolForMath @AccessDenied @SmoothMath @KingGeorge @TuringTest

OpenStudy (kinggeorge):

You're given a definite integral. Unless I'm severely mistaken, those should always converge for the same reason that the sum of a finite series is always finite.

OpenStudy (anonymous):

No. Definite integrals do not always converge....

OpenStudy (kinggeorge):

Interesting. May I have an example? (I'm assuming the function is defined and continuous on the interval as well)

OpenStudy (anonymous):

Consider: \[\frac{-\pi}{2}\int\limits_{0}^{\infty}(\frac{1}{2} - e^{-x})^{-1} dx \]

OpenStudy (kinggeorge):

In that case, I don't know how to make a convincing argument that this converges or diverges.